1:374 LTED Moodle Login Jacob MIII Flag question -1.5 1 in Section - 61 The maximum shear stress that will occur at section a how to get V
Answer #1

The Shear Force at Section a-a d=2in a *2221 Vm wc=50lb for the sections are Solb ► 2-AS Ala 9:40 = R² x 4R 3n NA (LED Zmar

Hope this helps

Please feel free to comment on any further queries or clarifications

All the best :)